OpenStudy (anonymous):

A machine gun is mounted on the top of tower 100m high.At what angle should the gun be inclined to cover a max range of firing the ground below?Muzzle speed=150m/s,take g=10m/s^2

OpenStudy (anonymous):

@siddhantsharan @RaphaelFilgueiras

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

range \[x=u _{x}t\cos (\theta)\]

OpenStudy (anonymous):

time taken for the journey \[u _{y}=u*\sin(\theta)\] \[=150(\sin \theta)\]

OpenStudy (anonymous):

Nop...@RaphaelFilgueiras ... U SEE..IT IS PROJECTED 100 M ABOVE...

OpenStudy (anonymous):

answer is 43.78'

OpenStudy (anonymous):

\[y=ut+\frac{ 1 }{ 2 }at^2\] \[100=150(\sin \theta)t+(1/2)10t^2\]

OpenStudy (anonymous):

use tan(theta)= ( 1+ 2gh/v^2)^(-1/2)

OpenStudy (anonymous):

now...how did i get this formula for maximum range... use equation of trajectory... y = h+ x tan(theta) - gx^2/(2v^2) sec^2(theta)

OpenStudy (anonymous):

for x= R(range), y=0

OpenStudy (anonymous):

form a quadratic equation in tan(theta)

OpenStudy (anonymous):

then use discriminant>=0 for real tan(theta)

OpenStudy (anonymous):

then u'll get a relation R<= v^2/g ( 1+ 2gh/v^2)^(1/2)

OpenStudy (anonymous):

now try by ur own...

OpenStudy (anonymous):

one more step... put the value of Rmax in the quadratic equation of tan(theta) then u will get the value of theta for which range will be maximum.

OpenStudy (anonymous):

angular manipulation not looking nice http://en.wikipedia.org/wiki/Range_of_a_projectile